value: 714 points Quad Enterprises is considering a new three-year expansion project that requires an initial fixed asset investment of $2.76 million. The fixed asset will be depreciated straight-line to zero over its three-year tax life, after which time it will be worthless. The project is estimated to generate $2,100,000 in annual sales with costs of $795,000. The tax rate is 34 percent and the required return on the project is 12 percent. What is the projects NPV? (Enter your answer in dollars, not millions of dollars, e.g. 1,234,567. Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.) NPV Hints ReferenceseBook & Resources Hint #1
Answer #1

Net present value = Present value of all the future cash flows - Initial cash outflow

Project life is of 3 years

Initial investment in fixed asset = $ 2.76 million

Depreciation on straight line basis = 2.76/3 = $ 0.92 million per year

Annual profit = Annual sales - costs - Depreciation

Annual profit = 2,100,000 - 795,000 - 920,000 = $ 385,000

Annual Cash flow = Annual profit - Taxes + Depreciation

Annual Cash flow = 385,000 - (385,000 * 34%) + 920,000

Annual Cash flow = 1,174,100

So, Net present value = 1,174,100*PVAF(12%,3) - 2,760,000

Net present value = 2,819,990.09 - 2,760,000

Net present value = $ 59,990.09